Which point lies on the LINE REPRESENTED BY THE EQUATION BELOW?
7x+6y=42
a)1,7
b)1,6
c)0,6
D)0,7

Answer by stanbon(75887)

Which point lies on the LINE REPRESENTED BY THE EQUATION BELOW?
7x+6y=42
Substitute each x/y pair into the equation to
see if you get a true statement:
Example:
a)(1,7)
Using (1,7): 7*1 + 6*7 = 42 ; 7+42 = 42 ; 49 = 42 ; not true
-----------------------------------------------------------------
b)(1,6)
c)(0,6)
D)(0,7)
Using (0,7): 7*0 + 6*7 = 42 : 0+42 = 42 ; 42 = 42 ; true
===========================================================
Cheers,
Stan H.
===============